OpenStudy (anonymous):

Solve 2/8x^2=1/8x I want to solve it using the LCD way but I don't know what to multiply both sides by.

OpenStudy (anonymous):

are the x's in the numerators?

OpenStudy (anonymous):

\[\frac{ 2 }{ 8x^2 } = \frac{ 1 }{ 8x }\] and no only in the denominators so I know that x can't equal 0.

OpenStudy (anonymous):

what about cross-multiplication?

OpenStudy (anonymous):

I'm not sure how to solve this by cross-multiplying...

OpenStudy (anonymous):

ok the LCM of the denominators is 8x^2 so multiply both sides by it

OpenStudy (anonymous):

\[8x^2\times \frac{ 2 }{ 8x^2 }=\frac{ 1 }{ 8x }\times8x^2\] So it'd look like this?

OpenStudy (anonymous):

yes now simplify

OpenStudy (anonymous):